OpenStudy (abb0t):

Challenge Algebra Question:

OpenStudy (abb0t):

SIMPLIFY: \[\frac{ (x^2+4)^\frac{ 1 }{ 2 }(2x)-x^2(\frac{ 1 }{ 2 })(x^2+4)^{-\frac{ 1 }{ 2 }}(2x) }{ x^2+4 }\]

OpenStudy (hba):

\[\frac{ 2 x \sqrt(x^2+4)-x^3/\sqrt{x^2+4} }{ x^2+4 }\] \[2x(x^2+4)-x^3/(x^2+4)^{3/2}\]

OpenStudy (hba):

\[\frac{ x^3+8x }{ (x^2+4)^{3/2}}\]

OpenStudy (abb0t):

Barnicles! Nicely done man!

OpenStudy (hba):

Thanks :D